A company has a policy of retiring company cars; this policy looks at number of miles driven, purpose of trips, style of car and other features. The distribution of the number of months in service for the fleet of cars is bell-shaped and has a mean of 62 months and a standard deviation of 4 months. Using the 68-95-99.7 rule, what is the approximate percentage of cars that remain in service between 50 and 58 months?
Answer #1

z score for 50 = (50 - 62)/4 = -3

z score for 58 = (58 - 62)/4 = -1

Hence by 68-95-99.7 rule:

Percentage = P(-3 < z < -1) = P(z < -1) - P(z < -3) = 16% - 0.15% = 15.85%